OpenStudy (anonymous):

Paul biked 14.6 miles on Saturday and 12.8 miles on Sunday. Mariah biked the same total distance as Paul, but only biked 11.7 miles on Saturday. How far did Mariah bike on Sunday? A. 15.7 miles B. 27.8 miles C. 13.5 miles D. 1.8 miles

OpenStudy (danjs):

They both traveled the same distance. What is the total distance that Paul traveled?

OpenStudy (anonymous):

idk

OpenStudy (danjs):

Paul biked 14.6 miles on Saturday and 12.8 miles on Sunday add the 2 numbers

OpenStudy (danjs):

Paul Biked total - 14.6 + 12.8 miles Maria biked - 11.7 + x which should be the same as paul

OpenStudy (anonymous):

just add the 2 numbers that Paul biked to get the total distance...since they both traveled the same distance. Then subtract the miles Mariah already biked and that should leave you with the answer.

OpenStudy (danjs):

14.6 + 12.8 = 11.7 + x solve for x

OpenStudy (superflytyguy):

Mariah needs to pick up the pace lol But @DanJS is right, 14.6 + 12.8 = 27.4 27.4 - 11.7 = 15.7

OpenStudy (superflytyguy):

A